OpenStudy (anonymous):

Use L'Hopital's rule to solve: (1-sinx)/(1+cos2x) LIM x --> pi/2

OpenStudy (ipwnbunnies):

What's the limit?

OpenStudy (anonymous):

sorry

OpenStudy (ipwnbunnies):

Cool. So L'Hopital's Rule only applies to indeterminite forms when you plug the limit in for x. In this case, we have a case of 0/0

OpenStudy (anonymous):

yes

OpenStudy (ipwnbunnies):

With L'Hopital's rule, we can still find the limit. Take the derivative of both the numerator and the denominator, and leave them in the numerator and denominator. Then, apply the limit again.

OpenStudy (anonymous):

i did. I got -cosx/-2sin2x

OpenStudy (anonymous):

this is 0/0?

OpenStudy (ipwnbunnies):

Sure. Apply it again! The fun never stops.

OpenStudy (anonymous):

i am getting sinx/-4Cos2x = -1/4

OpenStudy (anonymous):

@iPwnBunnies : the asked question has indeterminate form .. its 0/0 form

OpenStudy (ipwnbunnies):

Yes, the limit is -1/4. && I did address the form was 0/0. Did I miss something?

OpenStudy (ipwnbunnies):

Oh no, you're wrong! It's 1/4, not -1/4. Check your denominator math again lol.

OpenStudy (anonymous):

@iPwnBunnies : sorry i read it wrong..

myininaya (myininaya):

You could also just use straight algebra/trig if you preferred. Recall \[2\cos^2(x)=1+\cos(2x)\] so we could write... \[\frac{1-\sin(x)}{2\cos^2(x)}=\frac{1-\sin(x)}{2(1-\sin^2(x))}=\frac{1-\sin(x)}{2(1-\sin(x))(1+\sin(x))}\] Then those 1-sin(x) 's cancel \[\frac{1}{2(1+\sin(x))}\] now you can plug in the value it still get 1/4

OpenStudy (ipwnbunnies):

Trig subs are the worst. :|

OpenStudy (anonymous):

but how come if i use the other method I am getting a zero? - tht's what I don't get.\[\frac{ 1-sinx }{ 1+co2x }\] \[= \frac{ (1-sinx)(1+sinx )}{(1+ \cos2 )(1+sinx)}\] \[\frac{ 1-\sin ^{2}x }{ 1+sinxcos2x+sinx+\cos2x } = \frac{ \cos ^{2}x }{ 1+0+1+0 } =\frac{ 0 }{ 2 } =0\]

OpenStudy (anonymous):

or what's wrong with my method

myininaya (myininaya):

cos(2*pi/2)=cos(pi)=-1 not 0 so you have 1-1+1-1 on bottom not 1+0+1+0 look at the way I did it to get rid of the 0/0 case

OpenStudy (anonymous):

thanx @myininaya